OpenStudy (anonymous):

write an equation of the line containing the given point & parallel to the given line.....(6,-8) 8x-7y=2

OpenStudy (aroub):

parallel=same slope y=mx+b , m= 8/7 , (6,-8) -8=8/7(6)+b -8=48/7 +b -8-48/7=b -56/7-48/7=b -104/7=b so, y=8/7x -104/7

OpenStudy (anonymous):

Stupid COMPUTER!!!!!

OpenStudy (aroub):

lol